2log4-log3+2logx-4=0
Answer by Alan3354(21583)

2log4-log3+2logx-4=0
--------------
log(16) - log(3) + log(x^2) = 4
log((16x^2)/3) = 4
(16x^2)/3 = 10000
x^2 = 30000/16 = 1875
x =~ 43.301
Note: the negative root is not valid, since log(x) is used.